# Direct optimization of neoclassical ion transport in stellarator reactors

### v1

@@ 3.1 Configuration 1
Stage 1
Aspect Ratio 20 $10^{1}$
QP Symmetry Breaking 0 $10^{2}$
Stage 2
Aspect Ratio 20 $10^{1}$
QP Symmetry Breaking 0 $10^{1}$
$\epsilon_{\mathrm{eff}}$ $1\%$ $10^{2}$
Fuel-Ion-Relevant $D_{11}^{*}$ 0 $10^{2}$
Stage 3
$L_{11}^{e}$$\{-1.65\times 10^{-4},-5.16\times 10^{-4},-1.51\times 10^{-3}\}$ $10^{0}$
$L_{11}^{i}$ 0 $10^{0}$
$L_{31}^{e}$, $L_{31}^{i}$ 0 $10^{0}$
This optimization begins with a volume-averaged $\beta$ of roughly $3.5\%$. The $R_{00}$ Fourier component of the configuration is increased such that the initial aspect ratio is approximately 20. (Note that this increases the initial volume from the 1756 $\mathrm{m^{3}}$ quoted in section 2.) $R_{00}$ is then removed from the optimization space to keep the aspect ratio roughly constant.
Unlike the other presented configurations, Configuration 1 is optimized in several stages. In stage 1, quasipoloidal (QP) symmetry is targeted. This was found to bring the configuration to a more favorable region of optimization space (to a point), likely because lowering the value of the Boozer $B_{10}$ harmonic, the principle poloidal variation of $B$, tends to raise $L_{11}^{e}/L_{11}^{i}$ [31]. In stage 2, the weight on QP symmetry is lowered in favor of targets for $\epsilon_{\mathrm{eff}}$ and fuel-ion-relevant $D_{11}^{*}$. This is where the majority of the $D_{11}^{*}$ optimization occurs. In stage 3, SFINCS [51] is used to “polish” the magnetic configuration by directly targeting thermal transport coefficients using differential evolution implemented in SciPy [52]. The optimizer is allowed to alter only the Fourier harmonics used by VMEC to describe the plasma, except for $R_{00}$. SFINCS is directed to perform calculations for electrons and a deuterium-tritium “hybrid” species. Density and temperature profiles are chosen to be generically reactor-relevant and consistent with the choice of $\beta$ for the configuration. The radial electric field is arbitrarily set to $15~{}\mathrm{kV/m}$ on each of the three considered flux surfaces. The $L_{11}^{e}$ values from the end of stage 2 are used as targets in an attempt to hold radial electron transport roughly constant while driving $L_{11}^{i}$ (and therefore radial ion transport) to zero. The bootstrap current thermal transport coefficients $L_{31}^{e}$ and $L_{31}^{i}$ are also targeted and driven toward zero to minimize the bootstrap current of the final configuration. Details of the targets used for this configuration can be found in table 2.
Figure 1 shows the cross-sections for Configuration 1, which are highly elongated (as expected when QP optimization is used). The elongation creates some difficulties in performing neoclassical calculations for the outer flux surfaces since the Boozer spectra required to describe these flux surfaces are very large — the numerical values of neoclassical quantities are likely imprecise for $\rho\gtrsim 0.8$ in Configuration 1. Figure 2 shows that in the core, the fuel-ion-relevant $D_{11}^{*}$ are substantially reduced compared to W7-X high-mirror while the electron-relevant $D_{11}^{*}$ are somewhat elevated. The latter point is corroborated by the $\epsilon_{\mathrm{eff}}$ calculation shown in figure 3. These factors suggest this configuration will tend to have a relatively large $L_{11}^{e}/L_{11}^{i}$ in the core. For the NTSS simulation shown in figure 4, the reactor volume is scaled to $1450~{}\mathrm{m^{3}}$ and the on-axis magnetic field is scaled to $5~{}\mathrm{T}$. The simulation is initialized with an electron core temperature of $15.60~{}\mathrm{keV}$ and a fuel ion core temperature of $14.85~{}\mathrm{keV}$. The configuration has $L_{11}^{e}/L_{11}^{i}>1$ for $0\leq\rho\lesssim 0.4$ and consequently a strong electron root in this region. The bootstrap current is also small, and the configuration is coincidentally Mercier stable [53]. However, $\iota^{\prime}<0$ throughout, which makes the configuration vulnerable to neoclassical tearing modes [54], the aspect ratio is likely too large to be reactor-relevant, and the mirror ratio is very high (see table 1). Additionally, a final $P_{\alpha}$ of $600~{}\mathrm{MW}$ could not be achieved due to the high elongation of the configuration causing an errant divergence in the electric field near the last closed flux surface when the reactor volume was raised above 1450 $\mathrm{m^{3}}$. While Configuration 1 is not a fully developed reactor design, it demonstrates the conceptual effectiveness of separately optimizing neoclassical ion transport (through $D_{11}^{*}$) and electron transport (through $\epsilon_{\mathrm{eff}}$).
3.2 Configuration 2
Table 3: Optimization objectives used to produce Configuration 2 (section 3.2). All targets except aspect ratio are evaluated on the $\rho=\{0.25,0.5,0.75\}$ flux surfaces.
Objective Target Value Weight $\left(1/\sigma\right)$
Aspect Ratio 10.89 $10^{1}$
$\epsilon_{\mathrm{eff}}$ $1\%$ $10^{2}$
Fuel-Ion-Relevant $D_{11}^{*}$ 0 $10^{2}$
This optimization begins with a volume-averaged $\beta$ of roughly $4.3\%$. The $R_{00}$ Fourier harmonic is removed from the optimization space to prevent large changes in the aspect ratio through the course of the optimization. The philosophy of this optimization is to use a one-stage approach simpler than that of Configuration 1; consequently, only $\epsilon_{\mathrm{eff}}$, fuel-ion-relevant $D_{11}^{*}$, and aspect ratio are targeted, as can be seen in table 3. The first two targets are the most important and provide neoclassical optimization of the electrons and ions, respectively. The aspect ratio target is less important since $R_{00}$ is not included in the optimization space — it serves only to keep the aspect ratio in the neighborhood of the initial configuration by regulating relatively small fluctuations due to higher boundary modes.
From figure 1, Configuration 2 is significantly less elongated than Configuration 1, with cross-sections much more similar to scaled W7-X high-mirror. However, as seen in figure 2, the reduction of fuel-ion-relevant $D_{11}^{*}$ was even more substantial than in Configuration 1. From figure 3, $\epsilon_{\mathrm{eff}}$ was somewhat lower than in W7-X throughout the volume, staying below the targeted $1\%$ for $0\leq\rho\lesssim 0.85$. This results in lower electron-relevant $D_{11}^{*}$ values which compete with the small fuel-ion-relevant $D_{11}^{*}$ to make $L_{11}^{e}/L_{11}^{i}$ a bit smaller. For the NTSS calculations shown in figure 4, the reactor volume is scaled to $1900~{}\mathrm{m^{3}}$ and the on-axis magnetic field is scaled to $5~{}\mathrm{T}$. The initial core electron temperature is $24.39~{}\mathrm{keV}$ and the initial core fuel ion temperature is $23.21~{}\mathrm{keV}$. The figure shows that $L_{11}^{e}/L_{11}^{i}$ is only slightly larger than 1 for Configuration 2 in the region $0\leq\rho\lesssim 0.35$. An electron root appears in the same region with a peak $E_{r}$ value roughly half that of Configuration 1. The electric field profile for $\rho\gtrsim 0.35$ has a somewhat unconventional shape, likely because the $\epsilon_{\mathrm{eff}}$ profile is non-monotonic. While Configuration 2 has a favorable $E_{r}$ profile and a lower mirror ratio than Configuration 1, it is Mercier unstable, has $\iota^{\prime}<0$ for most flux surfaces, and has a relatively large bootstrap current. It should also be noted that, as shown in table 1, the input power required to achieve the NTSS results shown is higher than the final output power of the configuration — while the displayed profiles are self-consistent, steady-state solutions to the transport equations, the system “relaxes” into them from profiles requiring extremely high input power. This suggests that while it is possible to obtain an electron root by optimizing fuel ion transport such that the ions (rather than electrons) are the rate-determining species for the radial neoclassical flux, it is very challenging to do so. The challenge may be attributed to the $\sqrt{m_{s}}$ factor in equation 2.
3.3 Configuration 3
Table 4: Optimization objectives used to produce Configuration 3 (section 3.3). The neoclassical targets are evaluated on the $\rho=\{0.25,0.5,0.75\}$ flux surfaces. The rotational transform is evaluated on the $s=\rho^{2}=\{0,0.5,1\}$ flux surfaces.
Objective Target Value Weight $\left(1/\sigma\right)$
Aspect Ratio 10.89 $4\times 10^{-2}$
Average Boundary Elongation 1.87 $10^{-1}$
Volume $1755.76~{}\mathrm{m^{3}}$ $10^{-2}$
@@ 3.3 Configuration 3
This optimization begins with a volume-averaged $\beta$ of roughly $5.0\%$. Unlike Configurations 1 and 2, $R_{00}$ is included in the optimization space to allow for (small) changes to the aspect ratio, which is again one of the optimization targets. Volume and boundary elongation targets are also included in an attempt to prevent strong shaping like that found in Configuration 1. Values of these parameters from the initial configuration are used as targets. Rotational transform targets from the initial configuration are also included to help keep $\iota^{\prime}>0$ as this helps simultaneously improve MHD and neoclassical tearing mode stability. The same neoclassical targets were included as in section 3.2 and were the most important ingredients in the optimization. See table 4 for details.
@@
Figure 1 shows that the last closed flux surface of Configuration 3 is very similar to that of Configuration 2, and figure 2 shows similar $D_{11}^{*}$ optimization between Configurations 2 and 3. But unlike the other optimized configurations, Configuration 3 has an $\epsilon_{\mathrm{eff}}$ of almost exactly $1\%$ throughout the core. Because the fuel-ion-relevant $D_{11}^{*}$ are well-optimized and $\epsilon_{\mathrm{eff}}\approx 1\%$, we expect $L_{11}^{e}/L_{11}^{i}>1$ in the core. For the NTSS simulation, the reactor volume is scaled to $1900~{}\mathrm{m^{3}}$ and the on-axis magnetic field is scaled to $6~{}\mathrm{T}$. The initial core electron temperature is $21.46~{}\mathrm{keV}$ and the initial core fuel ion temperature is $18.57~{}\mathrm{keV}$. Figure 4 shows that $L_{11}^{e}/L_{11}^{i}$ is slightly greater than 1 for $0.15\lesssim\rho\lesssim 0.45$. Configuration 3 consequently has a weak ion root for $0\lesssim\rho\lesssim 0.05$ and an electron root for $0.05\lesssim\rho\lesssim 0.45$. The peak strength of $E_{r}$ in the electron root region is intermediate between that of Configurations 1 and 2. The shape of the electric field profile in the ion root region is similar to that of W7-X for the same values of $\rho$. The difficulty in achieving an electron root deep in the core (which was observed several times in our testing) can perhaps be explained by the fact that flux surface shapes always tend to become elliptical as $\rho\rightarrow 0$ [55]. Confinement (omnigenity) of electrons places constraints on the contours of $B_{\mathrm{min}}$, whereas confinement of fuel ions places constraints on the contours of $B_{\mathrm{max}}$ [56]. If cross-sections are almost purely elliptical, as is the case near the magnetic axis, it is difficult to optimize the magnetic field such that the latter constraints are (approximately) respected while the former are violated, which is necessary for the production of an electron root. Far from the axis, where there is more freedom to shape flux surfaces, this task becomes easier. In addition to an electron root in most of the core, Configuration 3 has a relatively small bootstrap current, nearly monotonically positive magnetic shear, and a reasonable power balance (see table 1). However, it is Mercier unstable. Again, we find that optimizing neoclassical ion and electron transport separately enables the production of a steady-state electron root at reactor-relevant collisionalities, but the task is challenging.

### v2

@@ 2 Formulation
Average Boundary Elongation 1.87 $10^{-1}$
Aspect Ratio 10.89 $4\times 10^{-2}$
Volume $1755.76~{}\mathrm{m^{3}}$ $10^{-2}$
$\chi^{2}=\sum_{i}\frac{\left(f_{i}-f_{i}^{\mathrm{target}}\right)^{2}}{\sigma_{i}^{2}},$ (3)
where $f_{i}$ is any quantity that is callable by the STELLOPT routines, $f_{i}^{\mathrm{target}}$ is the desired value for $f_{i}$, and $\sigma_{i}$ is a user-specified inverse weight. VMEC [37] is used as the MHD equilibrium solver in all cases. Differential evolution [38] and a Garabedian boundary representation [39] are utilized to find global minima more reliably. We allow STELLOPT to modify all the boundary modes of the initial configuration to minimize the objective function.
Figure 1: The pressure profile assumed during the finite-$\beta$ optimization of the example configuration.
The optimization begins from a fixed-boundary configuration based on the W7-X high-mirror configuration [40] whose volume and on-axis magnetic field strength have been scaled to the reactor-relevant values of roughly 1756 $\mathrm{m^{3}}$ and 5 T, respectively [41]. Stellarator symmetry and five field periods are assumed throughout. The initial value of $\beta$ is roughly $5\%$. The pressure profile used throughout the optimization can be seen in figure 1. Zero bootstrap current is assumed during the MHD calculations within the optimization loop. The core feature of the optimization is the minimization of an array of $D_{11}^{*}$ with fuel-ion-relevant $v_{E}^{*}$ and $\nu^{*}$. The $D_{11}^{*}$ values are calculated by DKES [42, 43], which is coupled to STELLOPT such that it can be called within the optimization loop. A $3\times 3$ array of $D_{11}^{*}$ values with $\left|E_{r}\right|/v=\left\{5.00\times 10^{-4},2.74\times 10^{-3},1.50\times 10^{-2}\right\}~{}\mathrm{T}$ and $\nu/v=\left\{2.50\times 10^{-6},1.58\times 10^{-5},1.00\times 10^{-4}\right\}~{}\mathrm{m^{-1}}$ is used. (These quantities, rather than their normalized counterparts, are specified as targets in STELLOPT.) The exact values of $\left|E_{r}\right|/v$ and $\nu/v$ seem relatively unimportant — the upper and lower values for each quantity are chosen to be approximately reactor-relevant bounds, and the middle values are the logarithmic means of the upper and lower values. Notably, modifications of $B_{0}$, $R_{0}$, and $\iota$ are relatively small in the optimization, so explicitly targeting $\left|E_{r}\right|/v$ and $\nu/v$ rather than $v_{E}^{*}$ and $\nu^{*}$ is acceptable. In addition to $D_{11}^{*}$, we target $\epsilon_{\mathrm{eff}}=1\%$ (calculated by NEO [1]) during the optimization. These neoclassical objectives are calculated for the $\rho=r/a=\{0.25,0.5,0.75\}$ flux surfaces. We also include objectives for the average boundary elongation, aspect ratio, and volume to prevent strong boundary shaping or large changes in the size of the configuration. Finally, we target the rotational transform on the $s=\rho^{2}=\{0,0.5,1\}$ flux surfaces in an attempt to keep $\iota^{\prime}>0$ as this simultaneously helps improve MHD [44] and neoclassical tearing mode [45] stability. Note that we do not include the full Mercier criterion as an optimization objective. The initial values of the MHD objectives are used as targets during the optimization — in other words, we seek to hold MHD properties steady during the optimization while modifying neoclassical properties. The neoclassical targets for $\epsilon_{\mathrm{eff}}$ and $D_{11}^{*}$ are the most important ingredients in the optimization, and the latter constitutes the novel approach of this work. See table 1 for numerical details of the objective function.
Once optimized, the transport characteristics of the example configuration are analyzed using DKES and NTSS [46]. NTSS performs flux-surface-averaged, time-dependent transport calculations (including momentum correction techniques [47]) to determine self-consistent profiles for density, temperature, radial electric field, bootstrap current, and other quantities on 51 radial grid points for a given equilibrium. It should be emphasized that the equilibrium itself is not updated during the transport calculations, but NTSS calculates bootstrap current using the $D_{31}^{*}$ database provided by DKES and updates the rotational transform profile by utilizing susceptance matrices [48]. NTSS determines the full profile of the radial electric field by solving a diffusion-like differential equation with hysteretic behavior [49] that minimizes the generalized heat production rate of the plasma [50]. This is necessary because the ambipolarity condition generally produces either one or three possible solutions for $E_{r}$ [21, 51, 52] and a discontinuous $E_{r}$ profile is nonphysical. In the case of three solutions, one is the ion root, one is the electron root, and one is an unstable root that is not physically realized. We choose the “diffusion coefficient” in the diffusion-like equation to be 2 $\mathrm{m^{2}s^{-1}}$ such that the transition region from the electron root to the ion root is roughly 5 cm wide, which is reasonable in light of global neoclassical simulations [33]. Specifying a detailed startup scenario is beyond our scope, so we initialize NTSS runs by instructing the software to “select” the electron root solution if it exists. This situation is relevant for reactors that primarily utilize electron heating for startup. When the configuration is loaded into NTSS, we rescale its volume to $1900~{}\mathrm{m^{3}}$ and its on-axis magnetic field strength to $6~{}\mathrm{T}$. This allows us to achieve a steady-state alpha power of roughly 600 MW, which approximately corresponds to a 3 GWth reactor. (We note that the electron root is expected to be more difficult to achieve as the magnetic field strength increases [3, 33], which may need to be taken into account if our methods are used alongside those for turbulence optimization.) Deuterium and tritium density profiles are fixed with on-axis values for each in excess of $10^{19}$ $\mathrm{m^{-3}}$ and core gradients relatively flat in $\rho$ — see figure 7 for details. We assume there are no heavy impurities in the plasma. The temperature profiles for each species are initialized with roughly parabolic shapes in $\rho$. The initial core electron temperature is set to $21.46~{}\mathrm{keV}$ and the initial core fuel ion temperature is set to $18.57~{}\mathrm{keV}$. NTSS uses the D-T cross-section to calculate the birth rate of alpha particles, which for simplicity are assumed to slow down on the flux surface of their birth. This determines the helium density, and quasi-neutrality then determines electron density. A very simple turbulent transport model is included based on fits to W7-AS data. It takes the form [53]
$Q_{s}^{\mathrm{turb}}=-\chi_{\mathrm{turb}}n_{s}T^{\prime}_{s},$ (4)
$\Gamma_{s}^{\mathrm{turb}}=-D_{\mathrm{turb}}n^{\prime}_{s},$ (5)
where $Q_{s}^{\mathrm{turb}}$ is the radial energy flux of species $s$ due to turbulence. Similarly to reference [29], we set $\chi_{\mathrm{turb}}=0.0065(P/\mathrm{MW})^{3/4}/\left(n_{s}/10^{20}~{}\mathrm{m^{-3}}\right)~{}\mathrm{m^{2}s^{-1}}$ and $D_{\mathrm{turb}}=0.0003(P/\mathrm{MW})^{3/4}/\left(n_{s}/10^{20}~{}\mathrm{m^{-3}}\right)~{}\mathrm{m^{2}s^{-1}}$ where $P=P_{\alpha}-P_{\mathrm{Br}}$, $P_{\alpha}$ is the power deposited by fast ions, and $P_{\mathrm{Br}}$ is the power radiated due to Bremsstrahlung. No external heating is considered. NTSS steps forward in time until it converges on a steady-state, self-consistent solution for the plasma profiles. In this work, we seek a configuration that can realize an electron root and maintain it during steady-state operation. (Configurations with insufficiently optimized ion confinement, such as scaled W7-X high-mirror, may be artificially seeded with an electron root during high-density operation but will quickly transition to an ion root.)
3 Results
Figure 2: Toroidal cross-sections of the last closed flux surface equally spaced in the toroidal angle across one half-field-period for (2(a)) the example configuration and (2(b)) scaled W7-X high-mirror.
Figure 3: Scans of $D_{11}^{*}$ versus $\nu^{*}$ with various $v_{E}^{*}$ at $\rho\approx 0.35$ for (3(a)) the example configuration and (3(b)) W7-X high-mirror. The upper and lower variational bounds on the $D_{11}^{*}$ values are calculated by DKES and bracketed by the “$+$” symbols, but these “error bars” are too small to see in most cases. The dotted lines indicate $D_{11}^{*}$ for a tokamak with the same aspect ratio, rotational transform, and (average) elongation as the given configuration for the same $\rho$. Note that the light blue and purple curves, roughly corresponding to fuel-ion-relevant $v_{E}^{*}$, are far more depressed in the example configuration than in W7-X high-mirror.
Figure 4: The geometric factor $\epsilon_{\mathrm{eff}}$ for the example configuration and W7-X high-mirror. These calculations are performed on the optimized equilibria without taking effects from the NTSS modeling (such as modifications to bootstrap current, $\beta$, etc.) into account.
Figure 5: Magnetic field strength of the example configuration at mid-radius with an on-axis magnetic field strength of roughly $5.19~{}\mathrm{T}$. $\theta$ is the VMEC poloidal angle and $\zeta$ is the VMEC toroidal angle.
Figure 6: Contours of the second adiabatic invariant $J$ for the example configuration. This plot was produced using the methods of reference [54] with an on-axis magnetic field strength of roughly $5.19~{}\mathrm{T}$. $\theta$ is the VMEC poloidal angle.
Figure 7: NTSS results for the steady-state density (7(a)), temperature (7(b)), radial electric field (7(c)), $L_{11}^{e}/L_{11}^{i}$ (7(d)), bootstrap current (7(e)), and $\iota$ (7(f)) profiles of the example configuration.
Table 2: Plasma parameters of the example configuration. The mirror ratio does not change throughout a simulation because it is determined by the magnetic geometry alone. The “Initial $P_{\alpha}$” and “Initial $P_{Br}$” entries refer to the alpha power and bremsstrahlung power, respectively, at the beginning of the NTSS simulation. All other entries refer to the end of the NTSS simulation. $\tau_{E}$ is the energy confinement time and $\tau_{E}^{\mathrm{ISS04}}$ is calculated using the ISS04 scaling expression.
Quantity Value
Mirror Ratio (%) 26.9
$\beta\right>_{\mathrm{Vol}}$ (%) 3.14
Initial $P_{\alpha}$ (MW) 495.6
Final $P_{\alpha}$ (MW) 601.5
Initial $P_{Br}$ (MW) 33.5
Final $P_{Br}$ (MW) 72.7
$\tau_{E}$ (s) $2.34=1.23\tau_{E}^{\mathrm{ISS04}}$
Figure 2 shows that the last closed flux surface of the example configuration is slightly more elongated than that of W7-X high-mirror, but thanks to the boundary elongation objective we employed, the difference is relatively small. Figure 3 shows that the optimization of fuel-ion-relevant $D_{11}^{*}$ for the example configuration has driven them to be substantially lower than those for W7-X high-mirror. The example configuration has also achieved its target $\epsilon_{\mathrm{eff}}=1\%$ almost perfectly throughout the core, as shown in figure 4. Because the fuel-ion-relevant $D_{11}^{*}$ are well-optimized and $\epsilon_{\mathrm{eff}}\approx 1\%$, we expect $L_{11}^{e}/L_{11}^{i}>1$ in the core. Additionally, figures 5 and 6 show that in the core, the example configuration is reasonably quasi-isodynamic and has contours of the second adiabatic invariant that decrease monotonically with increasing minor radius. Neither of these features are explicitly targeted during the optimization. The former feature may appear because the optimization drives the configuration toward having a magnetic field geometry that minimizes the radial drift of particles in the $\sqrt{\nu}$ regime (fuel ions), which is to say that the behavior of the fuel ions is nearly omigenous. The initial configuration employed for the optimization is approximately quasi-isodynamic, so as the optimizer searches for “nearby” field geometries with omnigenous fuel ion behavior, it is very likely to find another approximately quasi-isodynamic configuration. The latter feature may appear because omnigenity is necessary for the maximum-$J$ property to be present [55], so optimizing for omnigenous fuel ion behavior likely drove the configuration to be approximately maximum-$J$ by accident.
@@ 3 Results
Figure 7 shows that $L_{11}^{e}/L_{11}^{i}$ is slightly greater than 1 for $0.15\lesssim\rho\lesssim 0.45$. This configuration consequently has a weak ion root for $0\lesssim\rho\lesssim 0.05$ and an electron root for $0.05\lesssim\rho\lesssim 0.45$. The peak strength of $E_{r}$ in the electron root region is roughly $15~{}\mathrm{kV/m}$, which is slightly greater in magnitude than its peak strength in the ion root region. The shape of the electric field profile in the ion root region is similar to that of W7-X for the same values of $\rho$. The difficulty in achieving an electron root deep in the core (which was observed several times in our testing) can perhaps be explained by the fact that flux surface shapes always tend to become elliptical as $\rho\rightarrow 0$ [56]. Confinement (omnigenity) of electrons places constraints on the contours of $B_{\mathrm{min}}$, whereas confinement of fuel ions places constraints on the contours of $B_{\mathrm{max}}$ [57]. If cross-sections are almost purely elliptical, as is the case near the magnetic axis, it is difficult to optimize the magnetic field such that the latter constraints are (approximately) respected while the former are violated, which is necessary for the production of an electron root. There is more freedom to shape the flux surfaces far from the axis, so this task becomes easier. In addition to an electron root in most of the core, the example configuration has a relatively small bootstrap current (although we did not target this explicitly), almost monotonically positive magnetic shear, a reasonable power balance, and an energy confinement time $1.23$ times larger than that predicted by the ISS04 scaling expression [58] (see table 2). It is, however, Mercier unstable, which is unsurprising since Mercier stability was not among the optimization objectives.
4 Discussion
The previous example has shown that it is possible to separately optimize neoclassical electron and fuel ion transport by modifying stellarator magnetic geometry. Furthermore, this differential optimization may be exploited to obtain a steady-state electron root in the core of a stellarator reactor. However, we were unable to generate any completely satisfactory configurations despite substantial effort. The presented configuration, for instance, has a small ion root deep in the core and is Mercier unstable. Another configuration we generated used an optimization recipe very similar to the presented one to produce stronger fuel-ion-relevant $D_{11}^{*}$ suppression, but effects related to the slope of the $D_{11}^{*}$ vs $\nu^{*}$ curves (like those in figure 3) and the resonant electric field prevented an electron root from being realized. Another configuration was initially optimized for quasi-poloidal symmetry to encourage the development of the electron root by lowering the principle poloidal variation of $B$ [33], and while it had a very strong electron root throughout the core, was Mercier stable [44], and had very small bootstrap current, it was likely far too elongated to be reactor-relevant. Yet another configuration was produced by including only the aspect ratio and neoclassical targets from this work in its optimization recipe; it developed an electron root throughout the core but required too much initial heating power to be viable. These trials suggest that electron root optimization is challenging and that, while the fundamental idea of optimizing the electron and fuel ion neoclassical transport properties separately is sound, further development of the objective function would improve its reliability.
A possible extension of this work involves coupling a modern, fast neoclassical transport code such as KNOSOS [59, 60] to a modern optimization framework such as SIMSOPT [61] and performing optimizations similar to those presented in section 3. This would likely avoid some of the numerical difficulties we encountered while attempting to improve the optimization and allow for consideration of additional neoclassical effects. Notably, KNOSOS includes the tangential magnetic drifts (unlike DKES), which can be substantial for some configurations [62] and alter the confinement of fuel and impurity ions [63]. The tangential drifts can also make the electron root more difficult to achieve in maximum-$J$ configurations [64]. Because the maximum-$J$ property helps mitigate turbulence [65], it may be a useful optimization target for future devices. A code that takes the tangential drifts into account would therefore be helpful if an electron root is to be targeted as well. Due to its speed, KNOSOS can also calculate thermal transport coefficients very quickly. We performed some optimizations that used SFINCS [66] to target thermal transport coefficients as a “polishing” step after optimizing the monoenergetic transport coefficients. This was rather arduous but improved the neoclassical properties of the configurations, which suggests that targeting thermal transport coefficients throughout the optimization may be helpful — Lascas Neto and coworkers [67] have recently used SFINCS to do so and obtained promising results.
Another possible extension of this work involves developing a more theoretical approach to ensuring excellent neoclassical confinement of fuel ions and moderate confinement of electrons. Helander and coworkers [30] exploited the fact that in the formulation of omnigenity, there are separate conditions for zero average radial drift of marginally trapped particles (which pertains to fuel ions) and deeply trapped particles (which pertains to electrons). They derive an expression that allows the former conditions to be satisfied while the latter are violated and optimize a stellarator configuration using this expression as an objective. We suggest that further refinement of their analytical proxy, or perhaps a blend of the analytical and computational approaches, may provide the most reliable results for future stellarator optimization efforts.